OpenStudy (anonymous):

Juan is flying his kite on the football field. There is 70 meters of string between Juan and his kite. The string makes an angle of 38° with the ground. Find to the nearest meter how far above the ground the kite is flying. A) 43 meters B) 55 meters C) 89 meters D) 114 meters

OpenStudy (anonymous):

Let y be the how far above the ground the sky is flying... we know that \[\sin \theta= opposite side/hypothenuse\] in this case the opposite side is y and the hypothenuse is the string between Juan and his Kite \[\sin \theta=y/70 m\] \[70 m \times \sin 38=y\] \[43.096303272m=y\] Rounding off to the nearest meter we'll arrive at the answer y= 43 meters or simply letter A

OpenStudy (anonymous):

thank you !